Vapour pressure of an aquous solution is 2% less than that of the solvent the molality of the solution is

Grade:12th pass

1 Answers

Arun
6 years ago
P° - PS/ P° = 2/100 
m / m+ 55.5 = 2/ 100
 m / 55.5 m ≈ 1.13 m
 
hope it helps
if any doubt, let me know
thanks